# When Agents “Misremember” Collectively: Exploring the Mandela Effect in LLM-based Multi-Agent Systems

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced the capabilities of collaborative multi-agent systems, enabling them to address complex challenges. However, within these multi-agent systems, the susceptibility of agents to collective cognitive biases remains an underexplored issue. A compelling example is the Mandela effect, a phenomenon where groups collectively misremember past events as a result of false details reinforced through social influence and internalized misinformation. This vulnerability limits our understanding of memory bias in multi-agent systems and raises ethical concerns about the potential spread of misinformation. In this paper, we conduct a comprehensive study on the Mandela effect in LLM-based multi-agent systems, focusing on its existence, causing factors, and mitigation strategies. We propose ManBench, a novel benchmark designed to evaluate agent behaviors across four common task types that are susceptible to the Mandela effect, using five interaction protocols that vary in agent roles and memory timescales. We evaluate agents powered by several LLMs on ManBench to quantify the Mandela effect and analyze how different factors affect it. Moreover, we propose strategies to mitigate this effect, including prompt-level defenses (e.g., cognitive anchoring and source scrutiny) and model-level alignment-based defense, achieving an average 74.40% reduction in the Mandela effect compared to the baseline. Our findings provide valuable insights for developing more resilient and ethically aligned collaborative multi-agent systems. Our benchmark and code are available at https://github.com/bluedream02/Mandela-Effect.
@@ 1 Introduction
With the widespread deployment of large language models (LLMs) (Kojima et al., 2022), LLM-based multi-agent systems (Li et al., 2023a; Hong et al., 2024; Wu et al., 2025c) are increasingly used to address complex problems in fields like public policy analysis and social governance (Zhang et al., 2024). A key strength of these systems is their ability to simulate social dynamics like deliberation and consensus-building. However, this very capability introduces a significant risk: the emergence of collective cognitive biases analogous to those in human groups. For instance, the Mandela effect (Prasad and Bainbridge, 2022) is a well-known human cognitive bias where a group shares a false memory of a verifiable fact [^1] The Mandela effect, proposed by Fiona Broome in 2009, explains widespread false memories about South African anti-apartheid leader Nelson Mandela’s death, with many people incorrectly remembering that he died in prison in the 1980s, even though he actually passed away in 2013. Details of this effect are in Appendix B.. Just as this effect can arise from memory fragility and social reinforcement in people, LLM agents interacting within a system could develop similar distortions, leading to flawed group judgments. This could impact their collective problem-solving abilities or even raise significant ethical concerns. For example, the replication of collective cognitive biases might lead to shared false memories spreading through LLM interactions, causing agents to ignore the truth when influenced by suspicious evidence. Such effect undermines the reliability of AI-driven information systems and increases misinformation risks in high-stakes scenarios, including contract review (Narendra et al., 2024), and historical fact-checking (Tang et al., 2024).
@@ 2 Related Work
Social influence in LLMs. LLMs exhibit susceptibility to social influence, displaying conformity (Weng et al., 2025; Zhu et al., 2025), debate (Du et al., 2023), and sycophancy (Perez et al., 2023) tendencies. Other work has investigated how LLMs can be persuaded by user arguments (Xu et al., 2024). However, existing work focuses on short-term, in-context compliance. We advance the field by introducing long-term memory solidification—measuring whether socially-induced false beliefs become internalized into stable memories, a defining characteristic of the Mandela effect.
@@
Misinformation and factual robustness in LLMs. Most research focuses on the factuality of individual LLMs. A primary focus is on mitigating hallucinations, where models generate non-factual information (Elazar et al., 2021; Ji et al., 2023; Chen et al., 2024b; Ju et al., 2024; Liu et al., 2025; Li et al., 2023b; 2025). Some involve external verification of the outputs with trusted knowledge bases for fact-checking (Tang et al., 2024). In LLM-based multi-agent systems, Huang et al. (2025a) investigates the robustness of systems against malicious agents that intentionally inject incorrect information to disrupt group consensus. However, these studies mainly treat factual errors as technical failures. Our work addresses socially-induced misinformation, where false memories are caused by persuasive social contexts.
@@ 3.2 Interaction Protocols
(i) Group Composition. A key factor behind the Mandela effect is social influence that forms and strengthens false memories. We compare two types of social influences. The Generic Group simulates simple social consensus, where undifferentiated agents take turns talking without assigned roles, providing specious evidence to reinforce the consensus. In contrast, the Role-based Group uses a more advanced, narrative-driven interaction with five specialized agents who deliver complementary and strategically distinct evidence—each tailored to a specific role—to construct a multi-faceted false reality (Xu et al., 2026). We design five specialized agents: the Error Conclusion Initiator, who initially presents the false answer; the Detail Support Provider, who adds fabricated yet plausible details; the Group Consensus Reinforcer, who creates social proof by agreeing with the specious evidence to give the illusion of majority consensus; the Authority Endorser, who acts as an expert, using academic jargon to legitimize false memories; and the Questioning Compromiser, who first expresses doubt but is eventually convinced by the group’s narrative.
@@ C.2 Model
Claude 3.5 Haiku (Anthropic, 2024) claude-3-5-haiku-20241022
Claude 4 Sonnet (Anthropic, 2025) claude-sonnet-4-20250514
Gemini 2.5 Flash (Comanici et al., 2025) gemini-2.5-flash
Gemini 2.5 Pro (Comanici et al., 2025) gemini-2.5-pro
Open-source models LLama3.1-8B (Meta, 2024) Llama-3.1-8B-Instruct
LLama3.3-70B (Meta, 2025) llama-3.3-70b
Deepseek3.1 (Liu et al., 2024) deepseek-v3.1
@@
Qwen3-235B (Yang et al., 2025) qwen3-235b-a22b
@@ C.3 Agent Interaction Sequences for Varying Group Sizes
To systematically study the impact of group size on the AI Mandela Effect, we designed a series of interaction sequences for groups ranging from N==1 to N==15 agents. The construction of these sequences follows a psychologically-grounded, two-phase approach designed to model the progressive establishment and reinforcement of a collective false memory. The role counts and the full interaction sequence for each group size from N==1 to N==15 are detailed in Table 7.
@@ F.1 Evaluation and Defense
We evaluate our defenses on a 1,000-question subset of MedMCQA, a challenging medical question-answering dataset. We apply prompt-level defense to the closed-source model (GPT-4o) and model-level defense to the open-source model (Llama 3.1-8B), calculating the reality shift rate to show defenses’ effectiveness.

### v2

# arXiv:2602.00428v2[cs.CL] 01 Mar 2026

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced the capabilities of collaborative multi-agent systems, enabling them to address complex challenges. However, within these multi-agent systems, the susceptibility of agents to collective cognitive biases remains an underexplored issue. A compelling example is the Mandela effect, a phenomenon where groups collectively misremember past events as a result of false details reinforced through social influence and internalized misinformation. This vulnerability limits our understanding of memory bias in multi-agent systems and raises ethical concerns about the potential spread of misinformation. In this paper, we conduct a comprehensive study on the Mandela effect in LLM-based multi-agent systems, focusing on its existence, causing factors, and mitigation strategies. We propose ManBench, a novel benchmark designed to evaluate agent behaviors across four common task types that are susceptible to the Mandela effect, using five interaction protocols that vary in agent roles and memory timescales. We evaluate agents powered by several LLMs on ManBench to quantify the Mandela effect and analyze how different factors affect it. Moreover, we propose strategies to mitigate this effect, including prompt-level defenses (e.g., cognitive anchoring and source scrutiny) and model-level alignment-based defense, achieving an average 74.40% reduction in the Mandela effect compared to the baseline. Our findings provide valuable insights for developing more resilient and ethically aligned collaborative multi-agent systems [^1] Code and dataset are available at https://github.com/bluedream02/Mandela-Effect..
@@ 1 Introduction
With the widespread deployment of large language models (LLMs) (Kojima et al., 2022), LLM-based multi-agent systems (Li et al., 2023a; Hong et al., 2024; Wu et al., 2025c) are increasingly used to address complex problems in fields like public policy analysis and social governance (Zhang et al., 2024; Liu et al., 2026). A key strength of these systems is their ability to simulate social dynamics like deliberation and consensus-building. However, this very capability introduces a significant risk: the emergence of collective cognitive biases analogous to those in human groups. For instance, the Mandela effect (Prasad and Bainbridge, 2022) is a well-known human cognitive bias (Liu et al., 2025a) where a group shares a false memory of a verifiable fact [^2] The Mandela effect, proposed by Fiona Broome in 2009, explains widespread false memories about South African anti-apartheid leader Nelson Mandela’s death, with many people incorrectly remembering that he died in prison in the 1980s, even though he actually passed away in 2013. Details of this effect are in Appendix B.. Just as this effect can arise from memory fragility and social reinforcement in people, LLM agents interacting within a system could develop similar distortions, leading to flawed group judgments. This could impact their collective problem-solving abilities or even raise significant ethical concerns. For example, the replication of collective cognitive biases might lead to shared false memories spreading through LLM interactions, causing agents to ignore the truth when influenced by suspicious evidence. Such effect undermines the reliability of AI-driven information systems and increases misinformation risks in high-stakes scenarios, including contract review (Narendra et al., 2024), and fact-checking (Tang et al., 2024).
@@ 2 Related Work
Social influence in LLMs. LLMs exhibit susceptibility to social influence, displaying conformity (Weng et al., 2025), debate (Du et al., 2023), and sycophancy (Perez et al., 2023; Li et al., 2025a) tendencies. Other work has investigated how LLMs can be persuaded by user arguments (Xu et al., 2024). However, existing work focuses on short-term, in-context compliance. We advance the field by introducing long-term memory solidification—measuring whether socially-induced false beliefs become internalized into stable memories, a defining characteristic of the Mandela effect.
@@
Misinformation and factual robustness in LLMs. Most research focuses on the factuality of individual LLMs. A primary focus is on mitigating hallucinations, where models generate non-factual information (Elazar et al., 2021; Ji et al., 2023; Chen et al., 2024b; Ju et al., 2024; Liu et al., 2025b; Li et al., 2023b; 2025b). Some involve external verification of the outputs with trusted knowledge bases for fact-checking (Tang et al., 2024). In LLM-based multi-agent systems, Huang et al. (2025a) investigates the robustness of systems against malicious agents that intentionally inject incorrect information to disrupt group consensus. However, these studies mainly treat factual errors as technical failures. Our work addresses socially-induced misinformation, where false memories are caused by persuasive social contexts.
@@ 3.2 Interaction Protocols
(i) Group Composition. A key factor behind the Mandela effect is social influence that forms and strengthens false memories. We compare two types of social influences. The Generic Group simulates simple social consensus, where undifferentiated agents take turns talking without assigned roles, providing specious evidence to reinforce the consensus. In contrast, the Role-based Group uses a more advanced, narrative-driven interaction with five specialized agents who deliver complementary and strategically distinct evidence—each tailored to a specific role—to construct a multi-faceted false reality. We design five specialized agents: the Error Conclusion Initiator, who initially presents the false answer; the Detail Support Provider, who adds fabricated yet plausible details; the Group Consensus Reinforcer, who creates social proof by agreeing with the specious evidence to give the illusion of majority consensus; the Authority Endorser, who acts as an expert, using academic jargon to legitimize false memories; and the Questioning Compromiser, who first expresses doubt but is eventually convinced by the group’s narrative.
@@ C.2 Model
claude-3-5-haiku-20241022 (Anthropic, 2024) Claude 3.5 Haiku
claude-sonnet-4-20250514 (Anthropic, 2025) Claude 4 Sonnet
gemini-2.5-flash (Comanici et al., 2025) Gemini 2.5 Flash
gemini-2.5-pro (Comanici et al., 2025) Gemini 2.5 Pro
Open-source models Llama-3.1-8B-Instruct (Meta, 2024) Llama3.1-8B
llama-3.3-70b (Meta, 2025) Llama3.3-70B
deepseek-v3.1 (Liu et al., 2024) Deepseek3.1
@@
qwen3-235b-a22b (Yang et al., 2025) Qwen3-235B
@@ C.3 Agent Interaction Sequences for Varying Group Sizes
We designed a series of interaction sequences for groups ranging from N==1 to N==15 agents (Xu et al., 2026). The construction of these sequences follows a psychologically-grounded approach designed to model the progressive establishment and reinforcement of a collective false memory. The role counts and the full interaction sequence for each group size are detailed in Table 7.
@@ F.1 Evaluation and Defense
We evaluate our defenses on a 1,000-question subset of MedMCQA (Pal et al., 2022), a challenging medical question-answering dataset. We apply prompt-level defense and model-level defense to calculate the reality shift rate to show the effectiveness of defenses.